OpenStudy (anonymous):

x + 2y + z = 7 3x - y + z = -12 4x + 3y - 2z = 9

OpenStudy (anonymous):

Using rref\[\left[ \left[\begin{matrix}1 & 2 & 1 & 7 \\ 3 & -1 & 1 & 12 \\ 4 & 3 & -2 & 9 \end{matrix}\right] \right]\] \[\left[\begin{matrix}1 & 0 & 0 & -2 \\0 & 1 & 0 & 5\\0 & 0 & 1& -1\end{matrix}\right]\] x = -2 y = 5 z = -1